separate the differential equation into variables

xydx + (x+1)dy=0


1
Expert's answer
2022-02-02T12:37:14-0500

"xydx + (x+1)dy=0\\
xydx =- (x+1)dy\\
\frac{xdx}{x+1}=-\frac{dy}{y}"

Let, "x+1=t\Rightarrow dx=dt"

"\frac{t-1}{t}dt=-\frac{dy}{y}"

"(1-\frac{1}{t})dt=" "-\frac{dy}{y}"

Integrating both sides, we get

"t-\ln t=-\ln y+c"

"x+1-\ln (x+1)=-\ln y+c"

"x-\ln (x+1)+\ln y+1=c"